OpenStudy (anonymous):

(Cos theta - cos theta)^2+(cos theta + cos theta)^2 A. Sin^2 theta B. 4cos^2 theta C. 8 D. Cos^2 theta

OpenStudy (anonymous):

B

OpenStudy (anonymous):

because (Cos theta - cos theta)^2+(cos theta + cos theta)^2\[(Cos \theta - \cos \theta)^2 =0\]

OpenStudy (anonymous):

\[+(\cos \theta + \cos \theta)^2=(2\cos \theta)^2 =4\cos^2 \theta\]